# On Sparsity of integral points in orbits with big pullbacks under iterates

### v1

@@ Abstract.
We prove new unconditional results of sparsity of integral points on orbits under many maps in arbitrary dimensions, generalizing theorems of Yasufuku(2015) and others. The main ingredients are new diophantine approximation tools.
@@ 1. Introduction
    There exists $c_{j}\in\mathbb{Q}$ such that $\sum_{j=1}^{q}c_{j}{D_{j}}-(l-n+1)(n+1)A$ is big and $A-c_{j}D_{j}$ is $\mathbb{Q}$-nef for all $j\leq m$
@@
This theorem recovers and generalizes Yasufuku’s results when his $D_{(nc)}^{m}$ is a union of lines and his theorem becomes unconditional due to Schmidt’s subspace theorem. Yasufuku’s theorem can be made quantitative with respect to bounding the largest number of iterations occurring in an integral iterate (even if we allow the base field to vary with bounded degree) and also uniform not depending on the initial point of the orbit. This was recently achieved by Yasufuku in [21] using new diophantine approximation ingredients. His uniformity result replacing points in a given number field with algebraic points of bounded degree provides a version of the univariate results of Gunther and Hindes [5] for $\mathbb{P}^{N}.$ Similarly, we prove a non-quantitative algebraic points analogue result for projective varieties assuming bigness of some iterates’ pullbacks of a base divisor as suggested in [21, Remark 8]. For dynamics with one map, we state that below
@@
(See Theorem 4.4) Let $X$ be a projective variety of dimension $n$, $S$ be a finite set of places, and $\phi$ be an endomorphism of $X$. Suppose there exist $m$, a nontrivial effective divisor $D$ and $A$ an ample divisor of $X$ such that
@@
If $X=\mathbb{P}^{2}$ and the $D_{j}$’s are lines in general position and $A\sim O_{\mathbb{P}^{2}}(1)$, then condition (ii) can be reduced to $q>15/2$ thanks to a sharp diophantine approximation result by Levin [12, Theorem 1.5] (See Corollary 4.6).
In Section 2 we discuss some preliminaries on heights. The diophantine approximation tools are stated in Section 3, and the main results are proven in Sections 4 for nef and big pullbacks using He and Ru’s theorem and Le’s and Levin’s theorem on algebraic points respectively.
@@ 3. Tools from diophantine approximation
The next one is a new Schmidt subspace type of theorem for algebraic points.
@@
holds for all points $P\in X(\overline{k})\setminus\bigcup_{i}\text{Supp }D_{i}$ satisfying $[k(P):k]\leq\delta$, where $C(m,n,\delta)$ is a positive constant given in [11, Page 229].
@@
This is [11, Theorem 2], which improves the bound on [12, Theorem 5.2] together with [19, Lemma 3.4]. ∎
@@
[12, Theorem 1.5] Let $S$ be a finite set of places of a number field $k$ let $L_{1},...,L_{q}\subset\mathbb{P}^{2}$ be lines over $k$, in general position. Let $\epsilon>0$. There there exists a finite union of lines $Z\subset X$ such that for all points $P\in\mathbb{P}^{2}(\overline{k})\setminus Z$ satisfying $[k(P):k]\leq 2$ we have
@@ 4.1. Sparsity of integral points in orbits
Here we present some new results in the context of the new diophantine approximation result of He and Ru.
@@
Therefore, applying the above with $Q=\eta(P)$, combining with 4.3, 4.4, 4.5, and the condition in 4.1, we obtain
@@
Lemma 3.1 applied $m$ times to each system of divisors $\{D_{ij}\}_{j=0,...,q_{i}}$ implies that there exists a union $Z_{1}$ of varieties such that $\eta(P)\in Z_{1}^{\prime}$, or it belongs to a set $Z_{2}$ of bounded height. Therefore, this argument shows that if $\phi(P)$ is in the set 4.1, then $\phi(P)=\psi_{i}(\eta(P))\in\psi_{i}(Z^{\prime}_{i})\cup\psi_{i}(Z_{1})\cup\psi_{i}(Z_{2})$, which concludes the proof. ∎
@@
    $\psi_{i}^{*}(D)=D_{i1}+D_{i2}+\cdots+D_{i,q_{i}}+D_{i}^{\prime}$ is a sum of effective divisors, $D_{i1},\ldots,D_{i,q_{i}}$ are in $l_{i}$-subgeneral position for some $l_{i}\geq n$, $\sum_{j=1}^{q_{i}}\frac{D_{ij}}{d_{ij}}-(l_{i}-n+1)(n+1)A$ is big for all $i$ and $D_{ij}\equiv d_{ij}A$ where $d_{ij}$ are positive integers for all $i,j$.
  • (ii)
@@
    $(\phi^{(m)})^{*}(D)=D_{1}+D_{2}+\cdots+D_{q}+D^{\prime}$ is a sum of effective divisors, $D_{1},\ldots,D_{q}$ are in $l$-subgeneral position for some $l\geq n$, $\sum_{j=1}^{q}\frac{D_{j}}{d_{j}}-(l-n+1)(n+1)A$ is big and $D_{j}\equiv d_{j}A$ $d_{j}$ are positive integers for all $j$.
@@
4.2. Algebraic points of bounded degree
@@
    $\sum_{j=1}^{q_{i}}\frac{D_{ij}}{d_{ij}}-C(l_{i},n,\delta)A$ is big and $D_{ij}\equiv d_{ij}A$ $d_{ij}$ are positive integers for all $i,j$.
@@
Let $\phi_{1},\ldots,\phi_{\ell}$ be endomorphisms of $\mathbb{P}^{2}$, and $\mathcal{F}=\langle\phi_{1},\ldots,\phi_{\ell}\rangle$ all def. over a number $k$. Suppose there exist $\psi_{1},\ldots,\psi_{m}\in\mathcal{F}$ and a nontrivial effective divisor $D$ such that

### v2

# On Sparsity of integral points in orbits and correspondences with big pullbacks under iterates

@@ Abstract.
We prove new unconditional results of sparsity of integral points on orbits under many maps and correspondences in arbitrary dimensions, generalizing theorems of Yasufuku(2015) and others. The main ingredients are new diophantine approximation tools and recent constructions for correspondences due to Ingram (2011).
@@ 1. Introduction
    There exists $c_{j}\in\mathbb{Q}$ such that $\sum_{j=1}^{q}c_{j}{D_{j}}-(l-n+1)(n+1)A$ is big and $A-c_{j}D_{j}$ is $\mathbb{Q}$-nef for all $1\leq j\leq q$
@@
This theorem recovers and generalizes Yasufuku’s results when his $D_{(nc)}^{m}$ is a union of lines and his theorem becomes unconditional due to Schmidt’s subspace theorem. Yasufuku’s theorem can be made quantitative with respect to bounding the largest number of iterations occurring in an integral iterate (even if we allow the base field to vary with bounded degree) and also uniform not depending on the initial point of the orbit. This was recently achieved by Yasufuku in [22] using new diophantine approximation ingredients. His uniformity result replacing points in a given number field with algebraic points of bounded degree provides a version of the univariate results of Gunther and Hindes [5] for $\mathbb{P}^{N}.$ Similarly, we prove a non-quantitative algebraic points analogue result for projective varieties assuming bigness of some iterates’ pullbacks of a base divisor as suggested in [22, Remark 8]. For dynamics with one map, we state that below
@@
(See Thm.’s 5.4 and 5.5) Let $X$ be a projective variety of dimension $n$, $S$ be a finite set of places, and $\phi$ be an endomorphism of $X$. Suppose there exist $m$, a nontrivial effective divisor $D$ and $A$ an ample divisor of $X$ such that
@@
is not Zariski dense. Moreover, if $X=\mathbb{P}^{2}$, the $D_{j}$’s are lines in general position and $A\sim O_{\mathbb{P}^{2}}(1)$, then condition (ii) can be reduced to $q>15/2$.
Lastly, we obtain the first versions of these results in the context of correspondences, which generalizes the scenario of deterministic dynamical systems. By $C$ be a correspondence on $X$ we mean a subvariety $C\subset X\times X$ such that the coordinate projections are both finite and surjective. We think of the points $X(k)$ as a set of vertices, with the points $C(k)$ defining directed edges between those vertices, which induces a graph whose union of paths generalizes the union of orbits in a classical dynamical system. In this context, we prove analogous sparsity results for similar generalized integrality sets using constructions on correspondences that were made by Ingram [10] when there are divisors with large Correspondence Iterate Pullbacks (CIP). This is a concept that naturally generalizes pullbacks in dynamical systems and that we define in Section 4. In particular, for projective spaces, we obtain
Theorem 1.4.
(See Theorem 5.9) Let $C\subset\mathbb{P}^{N}\times\mathbb{P}^{N}$ be a subvariety (correspondence) of $\mathbb{P}^{N}$ defined over $k$, $S$ finite set of places of $k$. Let $\mathcal{P}$ be the scheme of paths associated to $C$, with associated maps $\pi:\mathcal{P}\rightarrow\mathbb{P}^{N}$ and $\sigma:\mathcal{P}\rightarrow\mathcal{P}$ as it is described in Section 4. Suppose that there exist $m\in\mathbb{N}^{>0}$ and a nontrivial effective divisor $D$ defined over $k$ that has an $m$-th iterate (CIP) $D^{(m)}$ with respect to $C$ such that
  • (i)
    $D^{(m)}=D_{1}+...+D_{q}+D^{(\text{rest})}$ is a sum of effective ample divisors where $D_{1},...,D_{q}$ are in $l$-general position for some $l\geq n$.
  • (ii)
    $\sum_{j=1}^{q}{\deg D_{j}}>C(l,n,\delta)$
Then for all sufficiently small $\epsilon$, the set
$\{\pi(\sigma^{(n)}(P)):P\in\mathcal{P},[k(\pi(\sigma^{(n)}(P))):k]\leq\delta,n\in\mathbb{N}^{\geq m},\sum_{\begin{subarray}{c}v\notin S\\ w\mid v\\ w\in M_{k(P)}\end{subarray}}\lambda_{w}(D,\pi(\sigma^{(n)}(P)))\leq\epsilon h(\pi(\sigma^{n}(P)))\}$
@@
Again, when $N=2$ and the $D_{j}$’s are lines in general position, we can replace $C(l,n,\delta)$ with the sharp 15/2 due to Levin[13], and the Zariski closed set containing the integrality set is a union of lines plus a finite set.
As another byproduct, we obtain the following novel univariate finiteness statement for integers in correspondences which in the particular setting of dynamical systems ( when $C$ is the graph of a rational function) agrees with the theorems of Silverman and Gunther-Hindes, holding for correspondences having an "iterate" with at least 3 poles,
Theorem 1.5.
(See Corollary 5.10) Let $C\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$ be a correspondence of $\mathbb{P}^{1}$ defined over a number field $k$. Let $\mathcal{P}$, $\pi:\mathcal{P}\rightarrow\mathbb{P}^{1}$ and $\sigma:\mathcal{P}\rightarrow\mathcal{P}$ as in Thm. 1.4. Suppose that the divisor at infinity ()($\infty$)( ∞ ) has an $m$-th iterate (CIP) $D^{(m)}$ that is effective and whose support has more than $C(1,1,\delta)$ points.
Then
$\{\pi(\sigma^{(n)}(P)):P\in\mathcal{P},[k(\pi(\sigma^{(n)}(P))):k]\leq\delta,n\in\mathbb{N}^{\geq m},\pi(\sigma^{(n)}(P))\text{ is an algebraic integer}\}$
is finite.
In particular, the subset of set above formed by the points defined over $k$ $(\delta=1)$ is finite when when we replace $C(1,1,\delta)$ with $2$
The theorem above is also true for $S$-integers with $S$ a finite set of places.
In Section 2 we discuss some preliminaries on heights and Section 4 contains the needed background on correspondences. The diophantine approximation tools are stated in Section 4, and the main results are proven in Sections 4,5 and 6, for nef divisors and pullbacks using He and Ru’s theorem, algebraic points and correspondences respectively.
@@ 3. Tools from diophantine approximation
The next one is a new Schmidt subspace type of theorem for algebraic points. The symbol $\equiv$ denotes numerical equivalence of divisors.
@@
holds for all points $P\in X(\overline{k})\setminus\bigcup_{i}\text{Supp }D_{i}$ satisfying $[k(P):k]\leq\delta$, where $C(m,n,\delta)$ is a positive constant given in [12, Page 229].
@@
This is [12, Theorem 2], which improves the bound on [13, Theorem 5.2] together with [20, Lemma 3.4]. ∎
@@
[13, Theorem 1.5] Let $S$ be a finite set of places of a number field $k$ let $L_{1},...,L_{q}\subset\mathbb{P}^{2}$ be lines over $k$, in general position. Let $\epsilon>0$. There there exists a finite union of lines $Z\subset X$ such that for all points $P\in\mathbb{P}^{2}(\overline{k})\setminus Z$ satisfying $[k(P):k]\leq 2$ we have
@@
4. Preliminaries on correspondences
Let $X$ be an algebraic variety defined over some field $k$, and let $C$ be a correspondence on $X$, by which we mean a subvariety $C\subset X\times X$ such that the coordinate projections are both finite and surjective.
We think of the points $X(k)$ as a set of vertices, with the points $C(k)$ defining directed edges between those vertices. There is a path space $\pi:\mathcal{P}\rightarrow X$ parametrizing paths through the aforementioned directed graph, with initial vertex indicated by $\pi$, as well as a shift map $\sigma:\mathcal{P}\rightarrow\mathcal{P}$ which corresponds to stepping forward once along the path (or, equivalently, forgetting the first vertex of a path). This is described in the first lemma below.
Let $x,y:C\rightarrow X$ be the coordinate projections. We will say that this correspondence is polarized if and only if there exists an ample Cartier divisor $L\in Div(X)\otimes\mathbb{R}$ and a real number $\alpha>1$ such that $y^{*}L$ is linearly equivalent to $\alpha x^{*}L$. It is easy to check that in the case where $C$ is the graph of a morphism $f:X\rightarrow X$, the correspondence is polarized just in case the algebraic dynamical system is polarized in the usual sense.
Lemma 4.1.
[10, Theorem 2.1] Let $X$ be a separated, integral $S$-scheme of finite type and let $C$ be a correspondence in the $S$-scheme $X$ (a subscheme of $X\times_{S}X$ such that the two projections are finite and surjective). Then there is a separated, integral, $S$-scheme $\mathcal{P}$, surjective morphisms $\pi:\mathcal{P}\rightarrow X,\epsilon:\mathcal{P}\rightarrow C$ and a finite surjective morphism $\sigma:\mathcal{P}\rightarrow\mathcal{P}$ such that $\pi=x\circ\epsilon$ and $\pi\circ\sigma=y\circ\epsilon.$
Now, we define a generalization of a pullback of an iterated map for the context of correspondences that will allow us to generalize integrality results to this setting.
Definition 4.2.
Let $D$ be a divisor of $X$ and $C$ a correspondence on $X$ with coordinate projections $x,y:C\rightarrow X$. We say that a divisor $D^{(n)}$ of $X$ is an $n$-th Correspondence Iterate Pullback (CIP) of $D$ with respect to $C$ if there are divisors $L_{1},...,L_{n-1}$ of $X$ such that, making $L_{0}=D$ and $L_{n}=D^{(n)}$, we have
$y^{*}L_{i-1}=x^{*}L_{i}\text{ for all }i=1,...,n.$
For our results, we will generally assume that an effective divisor $D$ of $X$ has an $m$-th (CIP) $D^{(m)}$ with respect to a correspondence $C$ such that $D^{(m)}=D_{1}+...+D_{q}+D^{\prime}$ is a sum of effective divisors where $D_{1},...,D_{q}$ are in subgeneral position.
@@ 5.1. Sparsity of integral points in orbits
Here we present some new results in light of the new diophantine approximation result of He and Ru.
@@
Therefore, applying the above with $Q=\eta(P)$, combining with 5.3, 5.4, 5.18, and the condition in 5.1, we obtain
@@
Lemma 3.1 applied $m$ times to each system of divisors $\{D_{ij}\}_{j=0,...,q_{i}}$ implies that there exists a union $Z_{1}$ of varieties such that $\eta(P)\in Z_{1}^{\prime}$, or it belongs to a set $Z_{2}$ of bounded height. Therefore, this argument shows that if $\phi(P)$ is in the set 5.1, then $\phi(P)=\psi_{i}(\eta(P))\in\psi_{i}(Z^{\prime}_{i})\cup\psi_{i}(Z_{1})\cup\psi_{i}(Z_{2})$, which concludes the proof. ∎
@@
    $\psi_{i}^{*}(D)=D_{i1}+D_{i2}+\cdots+D_{i,q_{i}}+D_{i}^{\prime}$ is a sum of effective divisors, $D_{i1},\ldots,D_{i,q_{i}}$ are in $l_{i}$-subgeneral position for some $l_{i}\geq n$,
  • (i)
    $\sum_{j=1}^{q_{i}}\frac{D_{ij}}{d_{ij}}-(l_{i}-n+1)(n+1)A$ is big for all $i$ and $D_{ij}\equiv d_{ij}A$ where $d_{ij}$ are positive integers for all $i,j$.
  • (iii)
@@
    $(\phi^{(m)})^{*}(D)=D_{1}+D_{2}+\cdots+D_{q}+D^{\prime}$ is a sum of effective divisors, $D_{1},\ldots,D_{q}$ are in $l$-subgeneral position for some $l\geq n$,
  • (ii)
    $\sum_{j=1}^{q}\frac{D_{j}}{d_{j}}-(l-n+1)(n+1)A$ is big and $D_{j}\equiv d_{j}A$ where $d_{j}$ are positive integers for all $j$.
@@
5.2. Algebraic points of bounded degree in orbits
@@
    $\sum_{j=1}^{q_{i}}\frac{D_{ij}}{d_{ij}}-C(l_{i},n,\delta)A$ is big and $D_{ij}\equiv d_{ij}A$ where $d_{ij}$ are positive integers for all $i,j$.
@@
Let $\phi_{1},\ldots,\phi_{\ell}$ be endomorphisms of $\mathbb{P}^{2}$, and $\mathcal{F}=\langle\phi_{1},\ldots,\phi_{\ell}\rangle$ all def. over a number $k$ with a finite set of places $S$. Suppose there exist $\psi_{1},\ldots,\psi_{m}\in\mathcal{F}$ and a nontrivial effective divisor $D$ such that